Please help me solve this word problem: An animal shelter spends $3.50 per day to care for each bird and $7.50 per day to care for each cat. Sarah noticed that the shelter spent $137.00 caring for birds and cats on Thursday. Sarah found a record showing that there were a total of 30 birds and cats on Thursday. How many birds were there on Thursday? How many cats were there on Thursday?
Answer by Cromlix(4381)

Hi there,
Make the cats = 'x'
and the birds = 'y'
x + y = 30.......(1)
350x + 750y = 13700......(20
Multiply (1) by 350
350x + 350y = 10500.....(1)
350x + 750y = 13700.....(2)
Subtract (2) from (1)
.......400y = 3200
..........y = 8
Substitute y = 8 into (1)
x + y = 30.......(1)
x + 8 = 30
x = 30 - 8
x = 22
22 birds and 8 cats
were in on Thursday.
Proof:
22($3.50) + 8($7.50) = $137.00
Hope this helps :-)